Have radiator that holds 16 quarts antifreeze. Currently has 16 quarts antifreeze 25% solution. How much antifreeze should we take out and then add back in to make a 40% solution of 16 quarts?
Answer by scott8148(6628)

let x="amount to be replaced"

(16-x)(25%)+(x)(100%)=(16)(40%) __ 4-.25x+x=6.4 __ .75x=2.4 __ x=3.2